OpenStudy (anonymous):

solve for b -1/6b =3 a.-18 B. -2 c. 2 d. 18

OpenStudy (anonymous):

Ok, since b is being multiplied by -1/6, you multiply both sides by -6/1, to get b = -18. The answer is a.

OpenStudy (anonymous):

thank you can you help me with another problem?

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

thank you its -3/7m =12 a. 28 b. 26 c.-28 d.-26

OpenStudy (anonymous):

Ok, so you multiply both sides by -7/3, so you get m = -28

OpenStudy (anonymous):

thank you so much!

OpenStudy (anonymous):

Your welcome!